# Minimax rates in variance and covariance changepoint testing

### v1

@@ 6.1 Auxiliary Lemmas
We first show that $\mathbb{P}(\mathcal{E}_{11}^{\complement})\leq\delta/4$. Without loss of generality, assume that $\lambda_{\max}^{s}(\Sigma_{1})\leq\lambda_{\max}^{s}(\Sigma_{2})$. We have
$\mathbb{P}\left(\mathcal{E}_{11}^{\complement}\right)$$\leq\mathbb{P}\left\{\widehat{\lambda}_{\max}^{s}\left(\widehat{\Sigma}_{1,\lceil\beta\rceil}\right)\geq c_{7}\lambda_{\max}^{s}(\Sigma_{1})\right\}.$ (291)
Now, Lemma 11 implies that
$\widehat{\lambda}_{\max}^{s}\left(\widehat{\Sigma}_{1,\lceil\beta\rceil}\right)$ $\leq s\left\lVert\widehat{\Sigma}_{1,\lceil\beta\rceil}\right\rVert_{\infty}.$ (292)
Let $Y_{j,k}$ denote the $(j,k)$-th element of $\widehat{\Sigma}_{1,\lceil\beta\rceil}$, for $j,k\in[p]$, so that
$Y_{j,k}=\frac{1}{\lceil\beta\rceil}\sum_{i=1}^{\lceil\beta\rceil}X_{i}(j)X_{i}(k).$ (293)
Due to Lemma 2.7.7 in Vershynin (2018) and Assumption 3, we have that
$\leq u\lambda_{\max}^{1}(\Sigma)$ (295)
$\leq u\lambda_{\max}^{s}(\Sigma),$ (296)
for all $i\in[n]$ and $j,k\in[p]$. For any such $j,k,$ we have
$\mathbb{P}\left\{|Y_{j,k}|>c_{7}\lambda_{\max}^{s}(\Sigma_{1})\right\}$$\leq\mathbb{P}\left\{|Y_{j,k}-\mathbb{E}Y_{j,k}|>c_{7}\lambda_{\max}^{s}(\Sigma_{1})-\mathbb{E}Y_{j,k}\right\}$ (297)
$\leq\mathbb{P}\left\{|Y_{j,k}-\mathbb{E}Y_{j,k}|>(c_{7}-1)\lambda_{\max}^{s}(\Sigma_{1})\right\},$ (298)
where the last inequality follows from the fact that $\mathbb{E}Y_{j,k}\leq\lambda_{\max}^{1}(\Sigma_{1})\leq\lambda_{\max}^{s}(\Sigma_{1})$.
From Bernstein’s Inequality (Theorem 2.8.1 in Vershynin 2018) we obtain
$\mathbb{P}\left[|Y_{j,k}-\mathbb{E}Y_{j,k}|\geq cu\lambda_{\max}^{s}(\Sigma_{1})\left(\sqrt{\frac{x}{\lceil\beta\rceil}}\vee\frac{x}{\lceil\beta\rceil}\right)\right]$ $\leq 2e^{-x},$ (299)
where $c\geq 1$ is the same absolute constant as in the proof of Lemma 2. Setting $x=\lceil\beta\rceil(c_{7}-1)/(cu)\geq\log(ep)(c_{7}-1)/c_{0}$, where $c_{0}=16cu$ is the constant from Lemma 2, we obtain
$\mathbb{P}\left\{|Y_{j,k}-\mathbb{E}Y_{j,k}|>(c_{7}-1)\lambda_{\max}^{s}(\Sigma_{1})\right\}$ $\leq 2{(ep)^{-\frac{c_{7}-1}{c_{0}}}}.$ (300)
By a union bound over all $j,k\in[p]$, it follows that
$\mathbb{P}\left(\mathcal{E}_{11}\right)$ $\leq\frac{2p^{2}}{(ep)^{(c_{7}-1)/c_{0}}}$ (301)
$\leq 2\exp\left(-\frac{c_{7}-1}{c_{0}}\right)$ (302)
$\leq\frac{\delta}{4},$ (303)
where we in the first inequality used that $(c_{7}-1)c_{0}\geq 2$ and that $c_{7}\geq c_{0}\log(8/\delta)$ in the second.
Now consider $\mathcal{E}_{12}$. Using the same arguments as in the proof of Lemma 7, we obtain
$\mathbb{P}(\mathcal{E}_{12}^{\complement})$ $\leq\delta/4,$ (304)
and the proof is complete. ∎
@@
Let $p\geq 1$, $n\geq 2$. For $i=1,2$, let $V_{i}\in\mathbb{R}^{pn\times pn}$ be given by $V_{i}=\text{Diag}(\left\{V_{j,i}\right\}_{j=1}^{n})$, where $V_{j,i}=\sigma^{2}I-\kappa_{i}u_{i}u_{i}^{\top}$ for $j\leq\Delta_{i}$ and $V_{j,i}=\sigma^{2}I$ for $j>\Delta_{i}$, where $\sigma>0$, $1\leq\Delta_{i}\leq n$, $0<\kappa_{i}<\sigma^{2}$ and $u_{i}\in\mathcal{S}^{p-1}$ for $i=1,2$. Assume that $\Delta_{1}\leq\Delta_{2}$. If $X\sim\text{N}(0,\sigma^{2}I)$, then
@@
$\Delta\left(\frac{\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}{\sigma^{2}-\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}\right)\wedge\left(\frac{\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}{\sigma^{2}-\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}\right)^{2}$ $\geq c\gamma,$ (321)
for some constant $c>0$. Then we have
@@
The following Lemma is a formalization of results in Section 5 in Bach et al. (2010).
@@
where $N(p,s)=\{Z\in\mathrm{PSD}(p)\ ;\ \mathrm{Tr}(Z)=1,\left\lVert Z\right\rVert_{1}\leq s\}$.
Note that the result in Bach et al. (2010) is slightly stronger as equalities prevail throughout in that paper. Since they do not prove the result, we prove the slightly weaker Lemma 11 here.
@@
Let $\Sigma\in\mathbb{R}^{p\times p}$ be a symmetric positive definite matrix. Then $\lambda_{\max}^{s}(\Sigma)$ is non-decreasing in $s$, and for any $s_{0}/2\leq s\leq s_{0}\leq p$ we have $\lambda_{\max}^{s_{0}}(\Sigma)\leq 4\lambda_{\max}^{s}(\Sigma)$. Moreover, if $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{p\times p}$ are two symmetric positive definite matrices, then $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})\leq\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})$ for any $s\in[p]$.
@@
Lemma 13 (Liu et al. 2021, Supplementary Material, Lemma 8).

### v2

@@ 6 Auxiliary Lemmas
We first show that $\mathbb{P}(\mathcal{E}_{11}^{\complement})\leq\delta/4$. Without loss of generality, assume that $\lambda_{\max}^{1}(\Sigma_{1})\leq\lambda_{\max}^{1}(\Sigma_{2})$. Observe that $\widehat{\lambda}_{\max}^{1}(A)=\lambda_{\max}^{1}(A)=\left\lVert A\right\rVert_{\infty}$ for any positive definite matrix $A$. In particular, we have
$\mathbb{P}(\mathcal{E}_{11}^{\complement})$$\leq\mathbb{P}\left\{{\lambda}^{1}_{\max}(\widehat{\Sigma}_{1,\lceil\log(ep)\rceil})>c_{3}\lambda_{\max}^{1}(\Sigma_{1})\right\}.$ (292)
Using the same arguments as in the proof of Lemma 6 it then follows that $\mathbb{P}(\mathcal{E}_{11}^{\complement})\leq\delta/4$. Moreover, using the same arguments as in the proof of Lemma 7, we obtain $\mathbb{P}(\mathcal{E}_{12}^{\complement})\leq\delta/4$, and the proof is complete. ∎
@@
Let $p\geq 1$ and $n\geq 2$. For $i=1,2$, let $V_{i}\in\mathbb{R}^{pn\times pn}$ be given by $V_{i}=\text{Diag}(\left\{V_{j,i}\right\}_{j=1}^{n})$, where $V_{j,i}=\sigma^{2}I-\kappa_{i}u_{i}u_{i}^{\top}$ for $j\leq\Delta_{i}$ and $V_{j,i}=\sigma^{2}I$ for $j>\Delta_{i}$, where $\sigma>0$, $1\leq\Delta_{i}\leq n$, $0<\kappa_{i}<\sigma^{2}$ and $u_{i}\in\mathcal{S}^{p-1}$ for $i=1,2$. Assume that $\Delta_{1}\leq\Delta_{2}$. If $X\sim\text{N}(0,\sigma^{2}I)$, then
@@
$\Delta\left\{\left(\frac{\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}{\sigma^{2}-\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}\right)\wedge\left(\frac{\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}{\sigma^{2}-\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})}\right)^{2}\right\}$ $\geq c\gamma,$ (309)
for some constant $c>0$, where $\lambda_{\max}^{s}(\cdot)$ is defined in (6). Then we have
@@
The following Lemma is a formalization of results in Section 5 in Bach et al., (2010).
@@
where $N(p,s)=\{Z\in\mathrm{PSD}(p)\ ;\ \mathrm{Tr}(Z)=1,\left\lVert Z\right\rVert_{1}\leq s\}$ and $\lambda_{\max}(\cdot)=\lambda_{\max}^{p}(\cdot)$ from (6).
Note that the result in Bach et al., (2010) is slightly stronger as equalities prevail throughout in that paper. Since they do not prove the result, we prove the slightly weaker Lemma 11 here.
@@
Let $\Sigma\in\mathbb{R}^{p\times p}$ be a symmetric positive definite matrix and let $\lambda_{\max}^{s}(\cdot)$ be defined as in (6). Then $\lambda_{\max}^{s}(\Sigma)$ is non-decreasing in $s$, and for any $s_{0}/2\leq s\leq s_{0}\leq p$ we have $\lambda_{\max}^{s_{0}}(\Sigma)\leq 4\lambda_{\max}^{s}(\Sigma)$. Moreover, if $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{p\times p}$ are two symmetric positive definite matrices, then $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})\leq\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})$ for any $s\in[p]$.
@@
Lemma 13 (Liu et al., 2021, Supplementary Material, Lemma 8).